The age of the father is twice the sum of the ages of his two children. After 20 years, his age will be equal to the sum of the ages of his children. What is the father's age?
Answer by ankor@dixie-net.com(22740)

The age of the father is twice the sum of the ages of his two children.
let f = father's age,
let c = sum of the children's age
f = 2c
:
After 20 years, his age will be equal to the sum of the ages of his children.
Both children will age 20 years, so their sum will be increased by 40 yrs
f + 20 = c + 40
f = c + 40 - 20
f = c + 20
Replace f with 2c
2c = c + 20
2c - c = 20
c = 20 is the sum of the children's age now
:
What is the father's age?
f = 2(20)
f = 40 yrs, father's age now